1.How many liters of 22% salt solution must be added to 75 liters of 67% salt to get a 47% salt solution?

2. How many gallons of a 40% alcohol solution should be mixed with 15 gallons of 80% solution to obtain a 60% alcohol solution?

Answer by ankor@dixie-net.com(22740)

1.How many liters of 22% salt solution must be added to 75 liters of 67% salt to get a 47% salt solution?
:
Let x = no. of liters of 22% solution:
:
Resulting amt = (x+75)
:
.22x + .67(75) = .47(x+75)
.22x + 50.25 = .47x + 35.25
.22x - .47x = 35.25 - 50.25
-.25x = -15
x = -15/-.25
x = +60 liters of 22% solution:
:
Check:
.22(60) + .67(75) = .47(60+75)
13.2 + 50.25 = .47(135)
:
:
2. How many gallons of a 40% alcohol solution should be mixed with 15 gallons of 80% solution to obtain a 60% alcohol solution?
:
Do this one exactly the same way. Email me if you need help on it.